OpenStudy (anonymous):

Chris's gas tank is 3/10 full. After he buys 5 gallons of gas, it is 1/2 full. How many gallons can Chris's tank hold?

OpenStudy (igreen):

3/10 + 5 = 1/2 Add 3/10 + 5

OpenStudy (igreen):

5 3/10 = 1/2 Multiply 2 * 5 3/10

OpenStudy (anonymous):

After that what do I do?

OpenStudy (igreen):

Just multiply 2 * 5 3/10, that will be your answer.

OpenStudy (anonymous):

3/10 *10?

OpenStudy (igreen):

No, that's: \((2) \times (5 \dfrac{3}{10})\)

OpenStudy (anonymous):

10.6

OpenStudy (igreen):

Yep, or \(10 \dfrac{3}{5}\)

OpenStudy (anonymous):

it says its incorrect

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@ilikemath

OpenStudy (igreen):

Oh, my bad..I thought it said it had 3/10 gallons in it now..

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

@iGreen